OpenStudy (anonymous):

Solve sin(arctan 3) ???

OpenStudy (amistre64):



OpenStudy (amistre64):

since the tangent is 3/1 you can construct a rt tri and solve for the hypot. then sin is just opp/hyp

OpenStudy (amistre64):

sin = 3/x where x = sqrt(1^2+3^2)

OpenStudy (anonymous):

X= ^10

OpenStudy (amistre64):

yep; x = sqrt(10) so sin = 3/sqrt(10) and simplify that to your hearts content

OpenStudy (anonymous):

What is the answer when you simplify?

OpenStudy (amistre64):

well, there are 2 ways to write this; and it depends on the author of the text you are using. some people abhor sqrts on the bottom so the multiply by sqrt/sqrt a/sqrt(n) = a sqrt(n)/n is a "simplified form". if they want a rounded number try to just input it into a calculator

OpenStudy (anonymous):

Ok how do you input it in a calculator?

OpenStudy (amistre64):

lol, depends on the calculator. id just open google and input: 3/sqrt(10) itll give you a calculator result